OpenStudy (anonymous):

write the prime factorization of the numerator and denominator. then write the fraction in the lowest terms....84/252

OpenStudy (anonymous):

so i need two answers i can try and draw if that will help

OpenStudy (anonymous):

84 = 2*2*3*7 252 = 2*2*7*3*3 1/3

OpenStudy (anonymous):

add and simplify.....22/66+2/66+8/66=

OpenStudy (anonymous):

wanna help with another?

OpenStudy (anonymous):

16/33

OpenStudy (anonymous):

between the pair of numbers insert the appropiate sign < = or >........3/5 ? 6/10

OpenStudy (anonymous):

=

OpenStudy (anonymous):

convert to mixed number....534/7

OpenStudy (anonymous):

76 2/7

OpenStudy (anonymous):

are 9/12 and 18/24 equivalent or not equivalent?

OpenStudy (anonymous):

9/12 = 3/4 18/24 = 3/4 equivalent

OpenStudy (anonymous):

24/40 type fraction in lowest terms using integer or fraction

OpenStudy (anonymous):

3/5

OpenStudy (anonymous):

36/5 divided by 9/10 in lowest terms as whole or mixed number if possible

OpenStudy (anonymous):

36/5 * 10/9 = 360/45 = 8

OpenStudy (anonymous):

i have more but it will take a few min to type them out or draw them

OpenStudy (anonymous):

I see, Im just goofing around, take yer time.

OpenStudy (anonymous):